# Poisoning Attacks and Defenses in Recommender Systems: A Survey

### v1

@@ 1 Introduction
Modern Recommender Systems (RS) have exhibited unprecedented capability in enhancing user experience by effectively matching the diverse preferences of individual users with an enormous number of available items Wang et al. (2023c); Cheng et al. (2016); Wang et al. (2019b). However, it has been increasingly recognized that these advanced RS show vulnerabilities to malicious activities, notably to Poisoning Attacks Gunes et al. (2014); Si and Li (2020); Zhang et al. (2020c). Such poisoning attacks integrate deceptive or harmful data into the RS training datasets or influence the model gradient propagation in the training process, to compromise system functionality and manipulate outcome recommendations. For instance, e-commerce RS may inadvertently promote low-quality products when an attacker manipulates a malicious user’s interactions with these products, thereby compromising the system’s integrity. Similarly, in the realm of news distribution, altering key phrases or visuals in news headlines or content can lead to the intentional dissemination of misinformation to specific user groups, potentially threatening societal stability. The presence of poisoning attacks undermines the original purpose of RS, transforming these platforms into tools that can be exploited for malicious intents.
@@
Initial research into poisoning attacks in RS primarily explored heuristic-based methods, yielding a wide variety of well-developed hand-engineered attack models. In its infancy stages, the studies focused primarily on basic attack models such as the primal sample attack, random attack, average attack Lam and Riedl (2004), and love/hate attack Turk and Bilge (2019), which necessitated minimal understanding of the rating data. As insights into attacker tactics deepened, researchers uncovered more sophisticated attacks. For example, the bandwagon attack Gunes et al. (2014) utilizes Zipf’s law of user-item relationships to strategically select items that strengthen connections between target items and popular items. The relation attack Yu et al. (2017) goes beyond mere item manipulation, targeting the dynamics within social recommendation algorithms. These methods involved pre-defining a fixed strategy for generating and injecting malicious profiles into the system. Nevertheless, the static nature of these strategies often led to their predictability, rendering them detectable to defense measures once their patterns are deciphered Wang et al. (2022d). Thus, there has been a paradigm shift towards more sophisticated and adaptable poisoning attacks. Recent studies in this field have increasingly concentrated on tailoring attack techniques to accommodate both decentralized and centralized RS architectures. Different architectures introduce distinct vulnerabilities, requiring attackers to develop more complex and customized strategies to exploit specific weaknesses within these systems Wang et al. (2022a); Yuan et al. (2023b); Yang et al. (2020a); Long et al. (2023). Additionally, a thorough understanding of the attackers’ capabilities within real-world scenarios is entailed  Wang et al. (2022d); Fan et al. (2021); Christakopoulou et al. (2019). In situations where the victim model or the details of the training dataset remain unknown, attackers are compelled to strategically allocate their economic resources to maintain stealth and effectiveness. These constraints represent substantial hurdles for attackers seeking to implement impactful strategies Lin et al. (2020); Zeng et al. (2023). Moreover, the choice of poisoning strategy—whether to manipulate data or to modify model gradients—is crucial. Each approach targets specific vulnerabilities within the RS, underscoring the necessity for strategic planning on the part of attackers in selecting the most effective method to achieve their intended impact.
@@
Gaining insights into the investigation of poisoning attacks, researchers have developed two primary defensive strategies for safeguarding RS: The first branch is to filter out poisoning information, which involves pinpointing and eliminating malicious elements within the system, such as counterfeit users exhibiting suspect behaviors or items infused with misleadingly negative information Xu and Zhang (2019); Wang et al. (2022b); Hao et al. (2023). This branch hinges on the application of supervised anomaly filtering, unsupervised clustering filtering, and confidence-based filtering strategies. The second strategy is to enhance the robustness of recommendation models. This approach includes bolstering model training, for example, through mutual enhancement among multiple models or incorporating adaptive learning of confidence levels for each sample into the training process Wu et al. (2021b); Wang et al. (2021, 2022c). It also involves information enrichment or augmentation, such as introducing additional data (e.g., attributes, text, images, social relationships) to enrich interaction data from existing data.
@@
  • We introduce a systematic pipeline that functions as a comprehensive taxonomy from the attacker’s perspective, aimed at providing a fresh understanding of poisoning attacks against recommendation systems.
@@
Paper Collection: In this survey, we conducted a thorough review of 48 latest high-quality research papers focusing on poisoning attacks and 46 research papers focusing on defense. Our literature search was conducted primarily using DBLP and Google Scholar, with the keywords ‘attack/poisoning + recommendation/recommender’ for poisoning attack research and keywords ‘defense/detection/robust + recommendation/recommender’. We then traversed the citations of the identified papers and included relevant studies. To capture the most influential and cutting-edge work, we consistently monitored leading conferences and journals, including IJCAI, ICDE, CIKM, KDD, WWW, SIGIR, WSDM, TKDE, etc. Additionally, to encompass emerging trends and innovative concepts, we included preprints from arXiv that presented novel ideas to this field. It should be noted that the overarching aim of this survey is to provide an updated landscape of this area. Consequently, this study delivers only a concise summary of those papers published before 2018 in this field instead of a detailed analysis.
@@ 2.1 Victim Recommender System
Let $\mathcal{U}$ and $\mathcal{I}$ respectively be the sets of users and items in RS, and $\mathcal{D}$ denotes the original user/item data, encompassing interaction information (e.g., clicks, likes, and collections) and features (e.g., text and images). A recommender $f_{\theta}$ is trained on dataset $\mathcal{D}$ to learn a low-dimensional representation set for predicting the potentially interested items from $\mathcal{I}$ to user $u\in\mathcal{U}$, formally:
@@
In the context of poisoning attacks, the recommendation model $f_{\theta}$ is the system that the attacker intends to manipulate, referred to as the ‘victim model’.
@@ 2.2 Attacks on Recommender Systems
In poisoning attacks, attackers can manipulate a set of malicious users $\mathcal{U}_{M}$. These users engage in strategic interactions with various items, occasionally involving modifications to some item features. The resultant malicious data $\mathcal{D}_{M}$ is then injected into the training data of RS. The intent behind such attacks is to manipulate the recommendation models’ learning process and ultimately achieve specific objectives. Given this context, we formulate the poisoning attack task as a bi-level optimization problem:
@@
$\mathrm{s.t.}$ $,\quad\mathbf{\Theta}^{*}=\arg\min\limits_{\mathbf{\Theta}}\mathcal{L}_{rec}(f_{\theta}(\mathcal{D},\mathcal{D}_{M})),$
where $\mathcal{L}_{attack}$ is the poisoning attack loss for evaluating attack utility, typically including measures such as the decrease in recommendation accuracy and the disruption in the ranking of specific target items. The poisonous data $\mathcal{D}_{M}$ are generated by alternately performing the inner optimizations for victim model training and outer optimizations. Specifically, the inner optimization focuses on training the victim model by considering both the original data and poisonous data. Meanwhile, the outer optimization is tasked with refining the poisonous data to achieve the intended attack objectives. Note that $f_{\theta}$ here can be a surrogate model as attackers might have no access to the victim model.
@@ 2.3 Defense against Poisoning Attacks
where $f_{filter}$ represents the malicious information filtering mechanism employed by the defender, where $\mathcal{D}^{\prime}$ is the high-quality training data obtained after filtering. $f_{robust}$ denotes the defender’s robustness enhancement strategy, resulting in $\mathbf{X}$, which consists of augmented information (e.g., enhanced representations, data confidence levels). Under these two defense strategies, the recommendation model leverages high-quality data and augmented information for training, thereby enhancing its resilience against poisoning attacks.
@@ 3.1.4 Implementing Poisoning Strategies
The architectural diversity of RS offers a playground for attackers, who can choose between two principal methods of infiltration: data poisoning and model poisoning, as depicted in Figure 5. In the strategy of data poisoning, an attacker inserts malicious interaction or multimodal data into the RS. This method involves contamination of the training dataset, where the injected data blends imperceptibly with legitimate data, subtly degrading the training process and skewing the model’s output in favor of the attacker’s goals Li et al. (2016); Fang et al. (2018). Model poisoning, in contrast, requires deeper, more technical manipulation. Here, the attacker targets the learning process itself by perturbing the model gradients. This involves the sophisticated generation and strategic upload of poisoned gradients during the aggregation phase, where information is pooled from multiple user devices devices Zhang et al. (2022a); Rong et al. (2022b). This tactic not only distorts the model’s learning trajectory but also evades detection by embedding within the collaborative learning process, making it a particularly stealthy and effective method of attack.
@@ 3.2.1 Poisoning Data Filtering for Defense
Poisoning data filtering is a crucial defensive strategy where defenders identify and remove elements such as fake users exhibiting suspicious behaviors, items infused with negative information, or unusual connections between users and items before the data is utilized within the system. We categorize this filtering defense into three distinct types: supervised anomaly filtering, unsupervised clustering filtering, and confidence-based filtering. Figure 7 illustrates the distinctions among these categories. Supervised anomaly filtering leverages labeled historical data to train models that can discern and isolate malicious entries based on learned patterns. This approach effectively creates a boundary in the vector space that distinguishes between normal and malicious information, thereby enabling the system to filter out harmful data proactively. Unsupervised clustering filtering is applied in scenarios lacking labeled data. It involves segregating data into various clusters without prior knowledge of their properties. This method identifies and filters out anomalies by detecting clusters that are significantly different in size or composition from the majority, isolating potentially malicious information based on its relational distance from normative clusters. Lastly, confidence-based filtering employs probabilistic or statistical models to assess and score each piece of data according to its likelihood of being malicious. This technique filters out data points with high anomaly scores, effectively removing entries most likely to compromise the system’s integrity.
@@ 3.2.2 Robust Training for Defense
Robust training employs strategies that inherently enhance the robustness of RS, allowing them to resist noise and attacks naturally. These strategies can be systematically categorized into three principal streams: collaborative enhancement training, confidence-aware training, and contextual enrichment training. Figure 8 illustrates the distinctions among these strategies. Collaborative enhancement training capitalizes on the synergy of multiple models working together, leveraging adversarial or collaborative methods to enhance each model’s resistance to attacks. This method not only strengthens individual model robustness but also fosters a cooperative defense mechanism across the system, enhancing overall resilience. Confidence-aware training, on the other hand, adaptively learns the malicious confidence level of each sample. By doing so, it prioritizes learning from reliable data and minimizes the impact of potentially poisoned inputs, thus maintaining the integrity and accuracy of the model’s outputs. The last stream is dedicated to the integration of supplementary data signals. Contextual enrichment training expands the training dataset with a wide variety of additional signals such as user attributes, textual content, images, and social relations. This diversity in data sources enriches the model’s contextual understanding, enabling it to make more informed and nuanced predictions.
@@ 4.1.1 System Degradation
In untargeted system degradation attacks, one common strategy is to select specific users or items and then predominantly poison the RS through these targeted samples. For instance, Infmix Wu et al. (2023a, 2021c) concentrates on identifying ‘influential users’ among malicious users who can exert a more significant impact on the system. The approach prioritizes the creation of interactions for these ’influential users’ to amplify the attack’s impact. Analogously, FedAttack Wu et al. (2022a) enhances the attack effect from the perspective of items by utilizing globally ’hardest items’ to disrupt model training. It selects items most and least relevant to user embeddings as the hardest negative and positive samples, respectively. Manipulating the interaction records of these hardest items significantly influences the system’s outcomes.
@@ 4.1.2 Targeted Manipulation
In a typical targeted attack scenario, attackers aim to significantly influence the placement of target items in the top-K recommendation lists for as many users as possible. To achieve this, attackers have designed sophisticated attack functions to optimize the attack model. For example, PoisonRecSong et al. (2021) focuses on maximizing the frequency of targeted items appearing in top-K recommendation lists by manipulating interaction data to boost item visibility across user profiles. Alternatively, A-hum Rong et al. (2022a) aims to directly maximize the predicted probability of each target item being recommended to users by influencing the model to assign higher relevance scores to these items during the recommendation process. Similarly, RAPU Zhang et al. (2021a) seeks to ensure that the predicted probabilities of the target item being recommended exceed those of other items by adjusting the training data to favor the target item, ensuring it consistently ranks higher in recommendation predictions. In addition to directly promoting or demoting specific items, some poisoning attack approaches design unique and specialized objectives to achieve similar outcomes. A notable example is GTA Wang et al. (2023b), which initially focuses on assessing user intent by predicting the most preferred item of each user and subsequently modifying the target item to resemble these preferred items closely. This operation updates the characteristics of the target item to align with the user-preferred items, subtly influencing the recommendation results.

### v2

@@ 1 Introduction
Modern Recommender Systems (RS) have exhibited unprecedented capability in enhancing user experience by effectively matching the diverse preferences of individual users with an enormous number of available items Wang et al. (2023c); Cheng et al. (2016); Wang et al. (2019b). However, it has been increasingly recognized that these advanced RS show vulnerabilities to malicious activities, notably to Poisoning Attacks Gunes et al. (2014); Si and Li (2020); Zhang et al. (2020c). Such poisoning attacks integrate deceptive or harmful data into the RS training datasets or influence the model gradient propagation in the training process, to compromise system functionality and manipulate outcome recommendations. For instance, e-commerce RS may inadvertently promote low-quality products when an attacker manipulates malicious user’s interactions with these products, thereby compromising the system’s integrity. Similarly, in the realm of news distribution, altering key phrases or visuals in news headlines or content can lead to the intentional dissemination of misinformation to specific user groups, potentially threatening societal stability. The presence of poisoning attacks undermines the original purpose of RS, transforming these platforms into tools that can be exploited for malicious intents.
@@
Initial research into poisoning attacks in RS primarily explored heuristic-based methods, yielding a wide variety of well-developed hand-engineered attack models. In its infancy stages, the studies focused primarily on basic attack models such as the primal sample attack, random attack, average attack Lam and Riedl (2004), and love/hate attack Turk and Bilge (2019), which necessitated minimal understanding of the rating data. As insights into attacker tactics deepened, researchers uncovered more sophisticated attacks. For example, the bandwagon attack Gunes et al. (2014) utilizes Zipf’s law of user-item relationships to strategically select items that strengthen connections between target items and popular items. The relation attack Yu et al. (2017) goes beyond mere item manipulation, targeting the dynamics within social recommendation algorithms. These methods involve pre-defining a fixed strategy for generating and injecting malicious profiles into the system. Nevertheless, the static nature of these strategies often led to their predictability, rendering them detectable to defense measures once their patterns are deciphered Wang et al. (2022d). Thus, there has been a paradigm shift towards more sophisticated poisoning attacks.
Recent studies in this field have increasingly concentrated on tailoring attack techniques to accommodate both decentralized and centralized RS architectures. Different architectures introduce distinct vulnerabilities, requiring attackers to develop more complex and customized strategies to exploit specific weaknesses within these systems Wang et al. (2022a); Yuan et al. (2023b); Yang et al. (2020a); Long et al. (2023). Additionally, a thorough understanding of the attackers’ capabilities within real-world scenarios is entailed  Wang et al. (2022d); Fan et al. (2021); Christakopoulou et al. (2019). In situations where the victim model or the details of the training dataset remain unknown, attackers are compelled to strategically allocate their economic resources to maintain stealth and effectiveness. These constraints represent substantial hurdles for attackers seeking to implement impactful strategies Lin et al. (2020); Zeng et al. (2023). Moreover, the choice of poisoning strategy—whether to manipulate data or to modify model gradients—is crucial. Each approach targets specific vulnerabilities within the RS, underscoring the necessity for strategic planning on the part of attackers in selecting the most effective method to achieve their intended impact.
@@
Gaining insights into the investigation of poisoning attacks, researchers have developed two primary defensive strategies for safeguarding RS: The first branch is to filter out poisoning information, which involves pinpointing and eliminating malicious elements within the system, such as counterfeit users exhibiting suspect behaviors or items infused with misleadingly negative information Xu and Zhang (2019); Wang et al. (2022b); Hao et al. (2023). This branch hinges on the application of supervised anomaly filtering, unsupervised clustering filtering, and confidence-based filtering strategies. The second strategy focuses on enhancing the robustness of recommendation models. This approach includes bolstering model training, for example, through mutual enhancement among multiple models or incorporating adaptive learning of confidence levels for each sample into the training process Wu et al. (2021b); Wang et al. (2021, 2022c). It also involves information enrichment or augmentation, such as introducing additional data (e.g., attributes, text, images) to enrich interaction data from existing data.
@@
  • We introduce a systematic pipeline that functions as a comprehensive taxonomy from the attacker’s perspective, aimed at providing a fresh understanding of poisoning attacks against recommender systems.
@@
Paper Collection: In this survey, we conducted a thorough review of 48 latest high-quality research papers focusing on poisoning attacks and 46 research papers focusing on defense. Our literature search was conducted primarily using DBLP and Google Scholar, with the keywords “attack/poisoning + recommendation/recommender” for poisoning attack research and keywords “ defense/detection/robust + recommendation/recommender”. We then traversed the citations of the identified papers and included relevant studies. To capture the most influential and cutting-edge work, we consistently monitored leading conferences and journals, including IJCAI, ICDE, CIKM, KDD, WWW, SIGIR, WSDM, TKDE, etc. Additionally, to encompass emerging trends and innovative concepts, we included preprints from arXiv that presented novel ideas to this field. It should be noted that the overarching aim of this survey is to provide an updated landscape of this area. Consequently, this study delivers only a concise summary of those papers published before 2018 in this field instead of a detailed analysis.
@@ 2.1 Victim Recommender System
Let $\mathcal{U}$ and $\mathcal{I}$ represent the sets of users and items in a RS, with $\mathcal{D}$ indicating the original user/item dataset that includes interaction details (such as clicks, likes, and collections) and features (like text and images). A recommender function $f_{\theta}$ is trained on the dataset $\mathcal{D}$ to learn a low-dimensional representation, which is used to predict items from $\mathcal{I}$ that might interest a user $u\in\mathcal{U}$. Formally:
@@
In the context of poisoning attacks, the recommendation model $f_{\theta}$ is the system that the attacker intends to manipulate, referred to as the “victim model”.
@@ 2.2 Attacks on Recommender Systems
Attackers can control a subset of malicious users $\mathcal{U}_{M}$, who engage in strategic interactions with various items, occasionally altering some item features. The resulting malicious dataset $\mathcal{D}_{M}$ is incorporated into the training data of the RS. The objective of these attacks is to distort the learning process of recommendation models to achieve specific goals. In this context, we define the poisoning attack task as a bi-level optimization problem:
@@
$\mathrm{s.t.}$ $,\quad\mathbf{\Theta}^{*}=\arg\min\limits_{\mathbf{\Theta}}\mathcal{L}_{rec}(f_{\theta}(\mathcal{D},\mathcal{D}_{M})).$
Here, $\mathcal{L}_{attack}$ represents the poisoning attack loss, which measures the attack’s effectiveness, typically considering metrics such as the decline in recommendation accuracy and the disruption of the ranking of targeted items. The malicious data $\mathcal{D}_{M}$ is generated through alternating inner and outer optimization steps. The inner optimization involves training the victim model using both the original and malicious data. The outer optimization refines the malicious data to fulfill the attack’s objectives. It’s important to note that $f_{\theta}$ can be a surrogate model, as attackers may not have access to the actual victim model.
@@ 2.3 Defense against Poisoning Attacks
where $f_{filter}$ represents the malicious information filtering mechanism employed by the defender, where $\mathcal{D}^{\prime}$ is the high-quality training data obtained after filtering. $f_{robust}$ denotes the defender’s robustness enhancement strategy, resulting in $\mathbf{X}$, which consists of augmented information (e.g., enhanced representations and data confidence levels). Under these two defense strategies, the recommendation model leverages high-quality data and augmented information for training, thereby enhancing its resilience against poisoning attacks.
@@ 3.1.4 Implementing Poisoning Strategies
The architectural diversity of RS offers a playground for attackers, who can choose between two principal methods of infiltration: data poisoning and model poisoning, as depicted in Figure 5. In the strategy of data poisoning, an attacker inserts malicious interaction or multimodal data into the RS. This method involves contamination of the training dataset, where the injected data blends imperceptibly with legitimate data, subtly degrading the training process and skewing the model’s output in favor of the attacker’s goals Li et al. (2016); Fang et al. (2018). Model poisoning, in contrast, requires deeper, more technical manipulation. Here, the attacker targets the learning process itself by perturbing the model gradients. This involves the sophisticated generation and strategic upload of poisoned gradients during the aggregation phase, where information is pooled from multiple user devices Zhang et al. (2022a); Rong et al. (2022b). This tactic not only distorts the model’s learning trajectory but also evades detection by embedding within the collaborative learning process, making it a particularly stealthy and effective method of attack.
@@ 3.2.1 Poisoning Data Filtering for Defense
Poisoning data filtering is a crucial defensive strategy where defenders identify and remove elements such as fake users exhibiting suspicious behaviors, items infused with negative information, or unusual connections between users and items before the data is utilized within the system. We categorize this filtering defense into three distinct types: supervised anomaly filtering, unsupervised clustering filtering, and confidence-based filtering. Figure 7 illustrates the distinctions among these categories. Supervised anomaly filtering leverages labeled historical data to train models that can discern and isolate malicious entries based on learned patterns. This kind of approaches effectively create a boundary in the vector space that distinguishes between normal and malicious information, thereby enabling the system to filter out harmful data proactivelyWang et al. (2022b); Hsiao et al. (2022). Unsupervised clustering filtering is applied in scenarios lacking labeled data. It involves segregating data into various clusters without prior knowledge of their properties. This kind of method identify and filters out anomalies by detecting clusters that are significantly different in size or composition from the majority, isolating potentially malicious information based on its relational distance from normative clusters Cai and Zhang (2019a, b). Lastly, confidence-based filtering employs probabilistic or statistical models to assess and score each piece of data according to its likelihood of being malicious. This technique filters out data points with high anomaly scores, effectively removing entries most likely to compromise the system’s integrity Liu (2020); Li et al. (2021).
@@ 3.2.2 Robust Training for Defense
Robust training employs strategies that inherently enhance the robustness of RS, allowing them to resist noise and attacks naturally. These strategies can be systematically categorized into three principal streams: collaborative enhancement training, confidence-aware training, and contextual enrichment training. Figure 8 illustrates the distinctions among these strategies. Collaborative enhancement training capitalizes on the synergy of multiple models working together, leveraging adversarial or collaborative methods to enhance each model’s resistance to attacks. This method not only strengthens individual model robustness but also fosters a cooperative defense mechanism across the system, enhancing overall resilience Wu et al. (2021b); Wang and Chen (2023). Confidence-aware training, on the other hand, adaptively learns the malicious confidence level of each sample. By doing so, it prioritizes learning from reliable data and minimizes the impact of potentially poisoned inputs, thereby preserving the integrity and accuracy of the model’s outputs Wang et al. (2021). The last stream is dedicated to the integration of supplementary data signals. Contextual enrichment training expands the training dataset with a wide variety of additional signals such as user attributes, textual content, images, and social relations. This diversity in data sources enriches the model’s contextual understanding, enabling it to make more informed and nuanced predictions Bian et al. (2021).
@@ 4.1.1 System Degradation
In untargeted system degradation attacks, one common strategy is to select specific users or items and then predominantly poison the RS through these targeted samples. For instance, Infmix Wu et al. (2023a, 2021c) concentrates on identifying “influential users” among malicious users who can exert a more significant impact on the system. The approach prioritizes the creation of interactions for these “influential users” to amplify the attack’s impact. Analogously, FedAttack Wu et al. (2022a) enhances the attack effect from the perspective of items by utilizing globally “hardest items” to disrupt model training. It selects items most and least relevant to user embeddings as the hardest negative and positive samples, respectively. Manipulating the interaction records of these hardest items significantly influences the system’s outcomes.
@@ 4.1.2 Targeted Manipulation
In a typical targeted attack scenario, attackers aim to significantly influence the placement of target items in the top-K recommendation lists for as many users as possible. To achieve this, attackers have designed sophisticated attack functions to optimize the attack model. For example, PoisonRecSong et al. (2021) focuses on maximizing the frequency of targeted items appearing in top-K recommendation lists by manipulating interaction data to boost item visibility across user profiles. Alternatively, A-hum Rong et al. (2022a) aims to directly maximize the predicted probability of each target item being recommended to users by influencing the model to assign higher relevance scores to these items during the recommendation process. Similarly, RAPU Zhang et al. (2021a) seeks to ensure that the predicted probabilities of the target item being recommended exceed those of other items by adjusting the training data to favor the target item, ensuring it consistently ranks higher in recommendation predictions. In addition to directly promoting or demoting specific items, some poisoning attack approaches design unique and specialized objectives to achieve similar outcomes. A notable example is GTA Wang et al. (2023b), which first evaluates user intent by predicting the most favored item for each user. It then adjusts the target item to closely mimic these preferred items. This process alters the target item’s attributes to match those of user-preferred items, subtly affecting the recommendation outcomes.